Some systems are under maintenance.

Watching

Watching

1 week ago —

No more issues have been observed since the suspected faulty optic has been disabled. A technician visit to swap the optic is pending. We will continue to monitor and intervene when necessary.

Investigating

1 week ago —

We have identified an potential failed optic, showing high amount of egress drops as root cause. The interface has been disabled and LACP is currently running on degraded redundancy. As this only impacts a small subset of total network capacity, enough redundancy is still in place.

1 week ago —

We are currently observing various brief disconnects in the length of ~30 seconds each for transport towards DRT campus. To avoid further impact, several carriers have been drained. We are awaiting feedback from partner network.

Incident UUID 203386fe-7789-41c0-89fe-678a50e929f5